I need to Write an equation of the line perpendicular to the line given by 2x + 3y = 7 and having the same y-intercept as the line given by 5x + 2y = 10. I know that I need to get the opposite reciprocal slope from the first equation and use the y-intercept from the second equation and put them together but I am still confused and need to be walked through what I need to do.

Answer by rfer(16322)

change the to slope intercept form
2x+3y=7
3y=-2x+7
y=-2/3x+7/3
y=-2/3x+2 1/3
------------------
5x+2y=10
2y=-5x+10
y=-5/2x+5
---------------
m=-2/3
b=5
-------------
can you go from here?